OpenStudy (anonymous):

derive the equation of the parabola with a focus at (0,-4) and a directrix of y=4

OpenStudy (anonymous):

f(x)=-16x2 f(x)=16x2 f(x)=-1/16x2 f(x)1/16x2

OpenStudy (dumbcow):

 a = -1/(4p) = -1/16 parabola is neg because it opens down